# D ECOR : Improving Coherence in L2 English Writing with a Novel Benchmark for Incoherence Detection, Reasoning, and Rewriting

### v1

@@ Abstract
Coherence in writing, an aspect that second-language (L2) English learners often struggle with, is crucial in assessing L2 English writing. Existing automated writing evaluation systems primarily use basic surface linguistic features to detect coherence in writing. However, little effort has been made to correct the detected incoherence, which could significantly benefit L2 language learners seeking to improve their writing. To bridge this gap, we introduce DECOR, a novel benchmark that includes expert annotations for detecting incoherence in L2 English writing, identifying the underlying reasons, and rewriting the incoherent sentences. To our knowledge, DECOR is the first coherence assessment dataset specifically designed for improving L2 English writing, featuring pairs of original incoherent sentences alongside their expert-rewritten counterparts. Additionally, we fine-tuned models to automatically detect and rewrite incoherence in student essays. We find that incorporating specific reasons for incoherence during fine-tuning consistently improves the quality of the rewrites, achieving a result that is favored in both automatic and human evaluations. [^1] We will make DECOR publicly available upon publication of the paper.
@@ 2.1 Definitions of coherence in English writing
2.2 Assessing coherence in texts
@@
Following the linguistic definition of coherence established in Reinhart (1980), a much recent work by Maimon and Tsarfaty (2023) incorporated these conditions into a novel benchmark, namely CoheSentia, and proposed a new coherence-annotation protocol that aligns better with human judgments. Unlike previous work that assigns a single holistic coherence score to each target text Lai and Tetreault (2018), CoheSentia provides incremental coherence labeling on a sentence-by-sentence basis, enabling humans to identify the specific reasons for incoherence. In our human annotation process, we follow the CoheSentia protocol to create the context-sentence pairs incrementally. We expand the linguistic fundamentals applied in CoheSentia and devise an annotation scheme that is tailored to incoherence detection and rewriting in L2 English writings.
@@ 3.1 Data Creation
We propose DECOR, a benchmark for assessing the writing coherence in L2 English essays. To construct the dataset, we first sampled $100$ medium-level essays from the TOEFL-11 dataset Blanchard et al. (2013), and corrected their basic grammar errors using the grammar models from Zhang et al. (2024). Then, we incrementally constructed context-sentence pairs $(C,S)$ for each essay, following the protocol suggested by Maimon and Tsarfaty (2023). In these pairs, sentence $S$ is the current sentence to be assessed, and context $C$ includes all preceding sentences in the essay up to and including the sentence immediately before $S$. Overall, we constructed $1,352$ $(C,S)$ pairs from the $100$ essays. The general statistics of DECOR is shown in Table 1.
@@ 3.2 Incoherence Detection Annotation Scheme
DECOR features the ability to detect the incoherence of a given context-sentence pair. Inspired by the linguistic fundamentals of coherence (i.e. cohesion, consistency, and relevance) defined in Reinhart (1980), we expanded these fundamentals with reference to previous work in order to apply the task of incoherence detection to L2 English writing. We describe five specific criteria for detecting incoherence in each context-sentence pair below.
@@
If the given context-sentence pair violates any of the aforementioned criteria, it is considered incoherent, necessitating the subsequent step (described in Section 3.3) to identify the specific reasons causing sentence $S$ to be incoherent to context $C$; otherwise, the sentence is labeled as coherent. The detailed annotation guidelines for this task are demonstrated in Appendix A.
@@ 4 Data annotation process and statistics
We calculated the inter-annotator agreement for these pairs using Cohen’s Kappa Cohen (1960). The two annotators achieved a $\kappa$ value of $0.83$ for Incoherence Detection, indicating an almost perfect agreement. For Incoherence Reasoning, they reached an average $\kappa=0.90$ across all reason types, also reflecting almost perfect agreement. The specific agreement scores for each reason type are presented in Appendix B. As for Incoherent Sentence Rewriting, the leading authors validated whether the new sentences are acceptable. In particular, a new sentence $S^{\prime}$ is acceptable if it preserves the semantic meaning of the original sentence $S$ and is coherent with the given context $C$. Overall, the rewrites by the two annotators were deemed acceptable at rates of 88% and 89%, respectively.
@@ 5.1.3 Data
Given the absence of a dedicated incoherence detection corpus for language learners suitable for model training purposes, we followed the approach recommended by Zhang et al. (2024) and synthesized task-specific incoherence detection data using GPT-4 OpenAI (2023). [^2] Throughout this paper, we employ GPT-4o as the default model unless otherwise specified. The prompt we used for GPT-4 is shown in Appendix D.1. To start with, we randomly sampled $800$ medium-level essays from the TOEFL-11 dataset and generated $11,267$ context-sentence pairs. We then used GPT-4 to analyze these pairs for incoherence, producing a label for each. In this process, $6,422$ sentences were identified as coherent, while $4,845$ were labeled as incoherent. For the training process, we allocated $90\%$ of this synthetic data for training purposes, denoted as $D_{T}$, and reserved the remaining $10\%$ for validation. Moreover, we also utilized out-of-distribution training data proposed in Maimon and Tsarfaty (2023), denoted as $D_{C}$.
@@ 5.1.4 Baselines
In light of the burgeoning field of powerful instruction-following models Ouyang et al. (2022); Bai et al. (2022); OpenAI (2023), we also explored two generation-based large language models: Llama 2 Touvron et al. (2023) and GPT-4. For Llama 2, we fine-tuned its 7B variant using our synthetic dataset $D_{T}$ for this task. With GPT-4, we tested in both zero-shot and 16-shot settings. Details of the prompts used in the GPT-4 experiments are provided in Appendix F.
@@ 5.1.5 Results and analysis
The results for the task of incoherence detection are demonstrated in Table 3. As observed, training with our task-specific synthetic dataset $D_{T}$ yielded superior results compared to using the out-of-domain dataset $D_{C}$. This improvement is attributed to the fact that $D_{C}$ consists solely of machine-generated texts, which introduces a significant distribution shift. Additionally, while GPT-4 with 16-shot examples surpassed all other models, smaller models trained on our synthetic data $D_{T}$, such as BERT-base and Llama-2-7B, achieved performance comparable to GPT-4 in a zero-shot setting. Moreover, DeBERTa-base matched GPT-4’s performance in the 16-shot setting and even exceeded it in the zero-shot scenario. We also experimented with combining both $D_{C}$ and $D_{T}$ during training; however, this did not lead to improved results. Details of the experiment are provided in Appendix E.
@@ 5.2.2 Evaluation metrics
In Figure 2(a), DECOR exhibits unbalanced annotations across the four reason types of incoherence. Hence, we report weighted F1 scores for each of the four sub-tasks to account for this imbalance.
@@ 5.2.3 Data
We adopted a similar approach as described in Section 5.1.3 to synthesize the training data for all four sub-tasks. Specifically, we prompted GPT-4 to identify all potential reasons for each instance of incoherence detected from Section 5.1.3, based on the seven predefined causes outlined in Table 2. The prompts we used for data synthesis are demonstrated in Appendix D.1. Furthermore, we post-processed the resulting data to create four distinct datasets, each serving as the training data for detecting Cohesion, Consistency, Relevance, and Others. For instance, in creating the training set for detecting Cohesion as the cause, an instance is labeled "Yes" if GPT-4 identifies R1, R2, or R3 as the cause of incoherence for that instance; otherwise, the label is "No", indicating that the incoherence is caused by other factors. Similar to 5.1.3, the synthetic datasets are denoted as $D_{T}$. The details for the post-processing and statistics of the resulting data for each sub-task are described in Appendix D.2.
@@ 5.2.4 Baselines
We adopted the same set of baseline models that are tested in the incoherence detection task: classification-based models (i.e. BERT and DeBERTa), and generation-based models (i.e. Llama 2 and GPT-4). Similarly, for each sub-task of the incoherence reasoning, each model predicts with two labels (i.e. yes or no) to determine if the sentence S is incoherent with the context C due to a specific cause. We fine-tuned BERT, DeBERTa, and Llama2-7B models on the task-specific synthetic data $D_{T}$ for each sub-task as well as the out-of-distribution data $D_{C}$. We also prompted GPT-4 under both zero-shot and 16-shot settings. The prompts for GPT-4 experiments are shown in Appendix F.
@@ 5.2.5 Results and analysis
The results for incoherence reasoning in terms of the four sub-tasks are demonstrated in Table 3. It was observed that training DeBERTa-base and Llama2-7B models with $D_{T}$ resulted in enhanced performance for Cohesion and Relevance when compared to training with $D_{C}$. For Cohesion, DeBERTa-base outperforms the Llama2-7B model and is close to the performance of GPT-4. In comparison, for the Consistency task, all of our models demonstrate markedly enhanced performance when trained with $D_{C}$ rather than $D_{T}$. This improvement is likely attributed to the imbalanced training data distribution in $D_{C}$, which more closely mirrors the Consistency class distribution in DECOR. For the task of Others, we have omitted $D_{C}$ from the table because the category Others is not included in $D_{C}$. Our Llama2-7B model, fine-tuned with $D_{T}$, achieved results comparable to GPT-4 in both zero-shot and 16-shot settings. We further explored the effects of combining $D_{T}$ and $D_{C}$ as training data to fine-tune our models for tasks excluding Others. The results varied across different tasks and are presented in Table 8 in Appendix E.
@@ 5.3.2 Evaluation metrics
We measured the systems’ performance of incoherence rewriting with the acceptance rate. This metric was determined by calculating the proportion of revised sentences $S^{\prime}$ that both achieve coherence with context $C$ and maintain minimally invasive edits, out of all evaluated incoherent context-sentence pairs. We specifically employed GPT-4 with 16-shot examples (with the best performance in the incoherence detection task) to determine if the rewrites $S^{\prime}$ are acceptable. Additionally, in line with the recent practices of evaluating instruction-following LLMs Zhou et al. (2024); Dubois et al. (2024), we asked GPT-4 to rank a pair of generated rewrites (one from the human-written reference, the other from the tested models) to decide which one is more coherent to the context $C$. For each tested model, we collect its win rate against the human reference. Note that we randomly shuffle the ordering of the pair-wise outputs to avoid position biases. The prompt we adopted for GPT-4 judging is shown in Appendix G.
@@ 5.3.3 Data
Given the reasons generated from the incoherence reasoning task, we prompted GPT-4 to generate the rewrites based on the identified reasons for incoherence. These rewrites are used as the training data for the incoherence rewriting task. The prompt we used for the rewrite synthesis and relevant statistics are shown in Appendix D.1. For automatic evaluation, we used all $213$ rewrites generated by our annotators, and we randomly selected a sample of $100$ for human evaluation.
@@ 5.3.5 Results and analysis
Moreover, we conducted a human evaluation where we asked our human expert to judge and compare system-generated rewrites with those produced by GPT-4.333To avoid biases, instead of the same annotators, we asked one of our leading authors to conduct the human evaluation. Additionally, the human evaluator was also tasked with a pairwise comparison between human-written references and the same set of GPT-4 rewrites. The results are shown in Figure 3. As expected, our human judges predominantly preferred rewrites produced by human experts over those generated by GPT-4, with the highest win rate reaching 74%. Consistent with the results in Table 4, fine-tuning Llama 2 with reasons for incoherence resulted in a higher win rate and a significantly lower loss rate compared to fine-tuning without reasons. A chi-square test indicates a significant difference between these two conditions (with p-value < 0.01). This supports our hypothesis that rewriting incoherent sentences with an understanding of their underlying causes produces higher-quality rewrites.
@@ 7 Limitations
Additionally, the texts sampled from the TOEFL-11 corpus for synthesizing our training data were limited by the specific writing prompts they addressed. This limitation may hinder the system’s ability to detect coherence in learner-produced writing that responds to out-of-domain prompts not included in the TOEFL-11 corpus. Future extensions of our work includes incorporating other L2 English writing corpus.
@@ 8 Ethics Statement
In this work, we utilized GPT-4 to synthesize our task-specific training data for coherence detection, reasoning, and rewriting. We also used it during the evaluation. To facilitate the reproducibility of our data synthesis process and evaluation results, we included all relevant prompts that were used in our paper. In addition, all the other models used in this research, are publicly available in peer-reviewed articles and referenced in this paper. All datasets, including our synthetic fine-tuning dataset and the annotated test set, will be released.
@@ A.2 Types of edits for incoherent sentence rewriting
Appendix B Inter-annotator agreement scores across different reason types

### v2

@@ Abstract
Coherence in writing, an aspect that second-language (L2) English learners often struggle with, is crucial in assessing L2 English writing. Existing automated writing evaluation systems primarily use basic surface linguistic features to detect coherence in writing. However, little effort has been made to correct the detected incoherence, which could significantly benefit L2 language learners seeking to improve their writing. To bridge this gap, we introduce DECOR, a novel benchmark that includes expert annotations for detecting incoherence in L2 English writing, identifying the underlying reasons, and rewriting the incoherent sentences. To our knowledge, DECOR is the first coherence assessment dataset specifically designed for improving L2 English writing, featuring pairs of original incoherent sentences alongside their expert-rewritten counterparts. Additionally, we fine-tuned models to automatically detect and rewrite incoherence in student essays. We find that incorporating specific reasons for incoherence during fine-tuning consistently improves the quality of the rewrites, achieving a result that is favored in both automatic and human evaluations. [^1] Data and code available: https://github.com/BillyZhang24kobe/writing2coherence
@@ 2.1 Definitions of coherence in English writing
2.2 Assessing coherence in machine-generated texts and human-written texts
@@
Following the linguistic definition of coherence established in Reinhart (1980), a more recent work by Maimon and Tsarfaty (2023) incorporated these conditions into a novel benchmark, namely CoheSentia, and proposed a new coherence-annotation protocol that aligns better with human judgments. Unlike previous work that assigns a single holistic coherence score to each target text Lai and Tetreault (2018), CoheSentia provides incremental coherence labeling on a sentence-by-sentence basis, enabling humans to identify the specific reasons for incoherence. In our human annotation process, we follow the CoheSentia protocol to create the context-sentence pairs incrementally. We expand the linguistic fundamentals applied in CoheSentia and devise an annotation scheme that is tailored to incoherence detection and rewriting in L2 English writings.
@@ 3.1 Data Creation
We propose DECOR, a benchmark for assessing the writing coherence in L2 English essays. To construct the dataset, we first sampled $100$ medium-level essays from the TOEFL-11 dataset Blanchard et al. (2013). Note that sentences from the TOEFL-11 dataset often have basic grammar mistakes like spelling errors or missing be-verbs, which can make the intended meaning unclear. Therefore, we used a grammar error correction model from Zhang et al. (2024) to fix these mistakes without changing the overall meaning of the sentence. Then, we incrementally constructed context-sentence pairs $(C,S)$ for each essay, following the protocol suggested by Maimon and Tsarfaty (2023). In these pairs, sentence $S$ is the current sentence to be assessed, and context $C$ includes all preceding sentences in the essay up to and including the sentence immediately before $S$. Overall, we constructed $1,352$ $(C,S)$ pairs from the $100$ essays. The general statistics of DECOR are shown in Table 1.
@@ 3.2 Incoherence Detection Annotation Scheme
Inspired by the linguistic fundamentals of coherence (i.e. cohesion, consistency, and relevance) defined in Reinhart (1980), we expanded these fundamentals with reference to previous work in order to apply the task of incoherence detection to L2 English writing. We describe five specific criteria for detecting incoherence in each context-sentence pair below.
@@
If the given context-sentence pair violates any of the aforementioned criteria, it is considered incoherent, necessitating the subsequent step (described in Section 3.3) to identify the specific reasons causing sentence $S$ to be incoherent to context $C$; otherwise, the sentence is labeled as coherent. The detailed annotation guidelines for this task are demonstrated in Appendix A. Note that annotators are instructed to evaluate the entire context when determining if a sentence is incoherent. If the context pertains to a fictional setting, such as a dream about fiction, these instances will not be considered incoherent.
@@ 4 Data annotation process and statistics
We calculated the inter-annotator agreement for these pairs using Cohen’s Kappa Cohen (1960). The two annotators achieved a $\kappa$ value of $0.83$ for Incoherence Detection, indicating an almost perfect agreement. For Incoherence Reasoning, they reached an average $\kappa=0.90$ across all reason types, also reflecting almost perfect agreement. The specific agreement scores for each reason type and more justifications for the annotation process are presented in Appendix B. As for Incoherent Sentence Rewriting, the leading authors validated whether the new sentences are acceptable. In particular, a new sentence $S^{\prime}$ is acceptable if it preserves the semantic meaning of the original sentence $S$ and is coherent with the given context $C$. Overall, the rewrites by the two annotators were deemed acceptable at rates of 88% and 89%, respectively.
@@ 5.1.3 Data
Given the absence of a dedicated incoherence detection corpus for language learners suitable for model training purposes, we followed the approach recommended by Zhang et al. (2024) and synthesized task-specific incoherence detection data using GPT-4 OpenAI (2023). [^2] Throughout this paper, we employ GPT-4o as the default model unless otherwise specified. The prompt we used for GPT-4 is shown in Appendix E.1. To start with, we randomly sampled $800$ medium-level essays from the TOEFL-11 dataset and generated $11,267$ context-sentence pairs. We then used GPT-4 to analyze these pairs for incoherence, producing a label for each. In this process, $6,422$ sentences were identified as coherent, while $4,845$ were labeled as incoherent. For the training process, we allocated $90\%$ of this synthetic data for training purposes, denoted as $D_{T}$, and reserved the remaining $10\%$ for validation. Moreover, we also utilized out-of-distribution (OOD) training data proposed in Maimon and Tsarfaty (2023), denoted as $D_{C}$. To our knowledge, $D_{C}$ is the only dataset featuring human annotations for coherence detection and incoherence reasoning in machine-generated texts. We incorporated $D_{C}$ as OOD training data to assess whether models trained exclusively on it could achieve strong performance in DECOR. This experiment aims to determine whether using OOD data can improve the detection of incoherence in texts authored by L2 speakers.
@@ 5.1.4 Baselines
In light of the burgeoning field of powerful instruction-following models Ouyang et al. (2022); Bai et al. (2022); OpenAI (2023), we also explored two generation-based large language models: Llama 2 Touvron et al. (2023) and GPT-4. For Llama 2, we fine-tuned its 7B variant using our synthetic dataset $D_{T}$ for this task. With GPT-4, we tested in both zero-shot and 16-shot settings. Details of the prompts used in the GPT-4 experiments are provided in Appendix G.
@@ 5.1.5 Results and analysis
The results for the task of incoherence detection are demonstrated in Table 3. As observed, training with our task-specific synthetic dataset $D_{T}$ yielded superior results compared to using the OOD dataset $D_{C}$. This improvement is attributed to the fact that $D_{C}$ consists solely of machine-generated texts, which introduces a significant distribution shift. Additionally, while GPT-4 with 16-shot examples surpassed all other models, smaller models trained on our synthetic data $D_{T}$, such as BERT-base and Llama-2-7B, achieved performance comparable to GPT-4 in a zero-shot setting. Moreover, DeBERTa-base matched GPT-4’s performance in the 16-shot setting and even exceeded it in the zero-shot scenario. We also experimented with combining both $D_{C}$ and $D_{T}$ during training; however, this did not lead to improved results. Details of the experiment are provided in Appendix F.
@@ 5.2.2 Evaluation metrics
In Figure 2(a), DECOR exhibits class imbalance across the four reason types of incoherence. Hence, we report weighted F1 scores for each of the four sub-tasks to account for this imbalance.
@@ 5.2.3 Data
We adopted a similar approach as described in Section 5.1.3 to synthesize the training data for all four sub-tasks. Specifically, we prompted GPT-4 to identify all potential reasons for each instance of incoherence detected from Section 5.1.3, based on the seven predefined causes outlined in Table 2. The prompts we used for data synthesis are demonstrated in Appendix E.1. Furthermore, we post-processed the resulting data to create four distinct datasets, each serving as the training data for detecting Cohesion, Consistency, Relevance, and Others. For instance, in creating the training set for detecting Cohesion as the cause, an instance is labeled "Yes" if GPT-4 identifies R1, R2, or R3 as the cause of incoherence for that instance; otherwise, the label is "No", indicating that the incoherence is caused by other factors. Similar to 5.1.3, the synthetic datasets are denoted as $D_{T}$. The details for the post-processing and statistics of the resulting data for each sub-task are described in Appendix E.2.
@@ 5.2.4 Baselines
We adopted the same set of baseline models that are tested in the incoherence detection task: classification-based models (i.e. BERT and DeBERTa), and generation-based models (i.e. Llama 2 and GPT-4). Similarly, for each sub-task of the incoherence reasoning, each model predicts with two labels (i.e. yes or no) to determine if the sentence S is incoherent with the context C due to a specific cause. We fine-tuned BERT, DeBERTa, and Llama2-7B models on the task-specific synthetic data $D_{T}$ for each sub-task as well as the out-of-distribution data $D_{C}$. We also prompted GPT-4 under both zero-shot and 16-shot settings. The prompts for GPT-4 experiments are shown in Appendix G.
@@ 5.2.5 Results and analysis
The results for incoherence reasoning in terms of the four sub-tasks are demonstrated in Table 3. It was observed that training DeBERTa-base and Llama2-7B models with $D_{T}$ resulted in enhanced performance for Cohesion and Relevance when compared to training with $D_{C}$. For Cohesion, DeBERTa-base outperforms the Llama2-7B model and is close to the performance of GPT-4. In comparison, for the Consistency task, all of our models demonstrate markedly enhanced performance when trained with $D_{C}$ rather than $D_{T}$. This improvement is likely attributed to the imbalanced training data distribution in $D_{C}$, which more closely mirrors the Consistency class distribution in DECOR. For the task of Others, we have omitted $D_{C}$ from the table because the category Others is not included in $D_{C}$. Our Llama2-7B model, fine-tuned with $D_{T}$, achieved results comparable to GPT-4 in both zero-shot and 16-shot settings. We further explored the effects of combining $D_{T}$ and $D_{C}$ as training data to fine-tune our models for tasks excluding Others. The results varied across different tasks and are presented in Table 8 in Appendix F.
@@ 5.3.2 Evaluation metrics
We measured the systems’ performance on the task of incoherence rewriting with the acceptance rate. This metric was determined by calculating the proportion of revised sentences $S^{\prime}$ that both achieve coherence with context $C$ and maintain minimally invasive edits, out of all evaluated incoherent context-sentence pairs. We specifically employed GPT-4 with 16-shot examples (with the best performance in the incoherence detection task) to determine if the rewrites $S^{\prime}$ are acceptable. Additionally, in line with the recent practices of evaluating instruction-following LLMs Zhou et al. (2024); Dubois et al. (2024), we asked GPT-4 to rank a pair of generated rewrites (one from the human-written reference, the other from the tested models) to decide which one is more coherent to the context $C$. For each tested model, we collect its win rate against the human reference. Note that we randomly shuffle the ordering of the pair-wise outputs to avoid position biases. The prompt we adopted for GPT-4 judging is shown in Appendix H.
@@ 5.3.3 Data
Given the reasons generated from the incoherence reasoning task, we prompted GPT-4 to generate the rewrites based on the identified reasons for incoherence. These rewrites are used as the training data for the incoherence rewriting task. The prompt we used for the rewrite synthesis and relevant statistics are shown in Appendix E.1. For automatic evaluation, we used all $213$ rewrites generated by our annotators, and we randomly selected a sample of $100$ for human evaluation.
@@ 5.3.5 Results and analysis
Moreover, we conducted a human evaluation where we asked our human expert to judge and compare system-generated rewrites with those produced by GPT-4.333To avoid biases, instead of the same annotators, we asked one of our leading authors to conduct the human evaluation. The detailed information for the human evaluation process is shown in Appendix D. Additionally, the human evaluator was also tasked with a pairwise comparison between human-written references and the same set of GPT-4 rewrites. The results are shown in Figure 3. As expected, our human judges predominantly preferred rewrites produced by human experts over those generated by GPT-4, with the highest win rate reaching 74%. Consistent with the results in Table 4, fine-tuning Llama 2 with reasons for incoherence resulted in a higher win rate and a significantly lower loss rate compared to fine-tuning without reasons. A chi-square test indicates a significant difference between these two conditions (with p-value < 0.01). This supports our hypothesis that rewriting incoherent sentences with an understanding of their underlying causes produces higher-quality rewrites.
@@ 7 Limitations
Additionally, the texts sampled from the TOEFL-11 corpus for synthesizing our training data were limited by the specific writing prompts they addressed. This limitation may hinder the system’s ability to detect coherence in learner-produced writing that responds to out-of-domain prompts not included in the TOEFL-11 corpus. Future extensions of our work include incorporating other L2 English writing corpora.
@@ 8 Ethics Statement
In this work, we utilized GPT-4 to synthesize our task-specific training data for coherence detection, reasoning, and rewriting. We also used it during the evaluation. To facilitate the reproducibility of our data synthesis process and evaluation results, we included all relevant prompts that were used in our paper. In addition, all the other models used in this research, are publicly available in peer-reviewed articles and referenced in this paper. All datasets, including our synthetic fine-tuning dataset and the annotated test set, are released.
@@
9 Acknowledgement
We would like to thank Yanda Chen, Ryan Shea, and people from the Columbia NLP group for their valuable discussions on the paper. We also thank all reviewers for their constructive feedback and suggestions, which significantly improve our work. In addition, we extend our gratitude to our expert annotators for their time and contributions to the completeness of the benchmark.
@@ A.2 Types of edits for incoherent sentence rewriting
Appendix B More details about inter-annotator agreement
B.1 Inter-annotator agreement scores across different reason types